# A systematic assessment of Large Language Models for constructing two-level fractional factorial designs

### v1

@@ 3.1.3 Zero-shot chain of thought
A chain of thought (CoT) is a series of intermediate natural language reasoning steps that can be added to a prompt to enhance the output of an LLM (wei2022chain). This technique allows the LLM to decompose the main problem into a series of smaller intermediate problems, each solved before giving the final output. wei2022chain show that adding examples written using CoT to prompts improves the performance of transformer-based LLMs in arithmetic, symbolic reasoning, and common sense tasks.
kojima2022large show that CoT can also improve the quality of zero-shot prompts. Specifically, they show that adding the sentence “Let’s think step by step,” or a similar text to a zero-shot prompt, generally enhances the performance of transformer-based LLMs in similar tasks as wei2022chain. kojima2022large refer to the technique of adding this sentence (or one of its variants) to a zero-shot prompt as Zero-shot-CoT. Attractive features of this technique include its simplicity, task independence, and the fact that it avoids the need for good CoT-based examples, which require significant expertise and careful writing.
@@ 3.1.4 Output format
We can set a format for an LLM to tailor the structure of their output to our needs (white2023prompt). For example, we can provide constraints on the length and format of its output in the prompt. An output format is particularly useful for our problem because the LLMs tend to provide two-level designs with additional text to explain its properties and construction; see ChatGPT’s output 1 and Gemini’s output A1 in the Aoppendix. However, for our numerical experiments, we only need a table with the design in a specific format. That is, the table must have $m$ columns and $n$ rows with entries equal to 1-1 or 1. The table must also have a header row with the labels of the factor columns and an additional column to label the runs.
@@
The second element of our output format allows us to streamline the LLMs output because it sets the format for the table. Specifically, it instructs the LLM to generate the table using a comma-separated values (CSV) format, which is common for storing tables, matrices, and arrays. In this format, every entry in the table is separated by a comma. We specify the end of a row using two backslash symbols (\\). To set our second format element, we add the following sentences to our prompt: “The table must be in a comma-separated values (CSV) format. Specifically, the values in the table must be separated by ‘,’ and each row must end with ‘\\’. In the table, the first row will be used as a header row to label the factors using the letters of the English alphabet starting with ‘A’. The first column will be called “Run” and used to count the number of runs starting at ‘1’. Each design cell (excluding the header and Run columns) must contain either ‘-1’ or ‘1’.” We instruct the LLMs to label factors using the letters of the English alphabet following montgomery2017design, ChatGPT’s output 1, and Gemini’s output A1. Moreover, the maximum number of factors in our experiments in Section 4 is 26, which is the number of letters in this alphabet.
@@ 3.2 Setup for the LLMs
To construct designs with Prompt 2, we use the GPT-5.1 and Gemini 2.5 Flash models, which are the default LLMs in ChatGPT’s and Gemini’s web interfaces, respectively, at the time of writing this article. Technically, these LLMs are gpt-5.1-chat-latest (openai_gpt5_1_chat_latest) and gemini-2.5-flash (google_gemini_2_5_flash). We run them using the Python packages called openai (openai_python) and google-genai (google-genai), and API keys from OpenAI and Google. An API key is a unique code that can be used to access and communicate with an API, which, in turn, provides access to an LLM. The openai and google-genai packages provide user-friendly Python functions to interact with the APIs.
@@
The Python packages and APIs enable more tuning parameters of the LLMs than ChatGPT’s and Gemini’s web interfaces. For example, through the google-genai package, we can set the parameters called temperature, top-k, and top-p of gemini-2.5-flash, which control the randomness and diversity of its output to multiple runs of the same prompt (vertex-ai_content_generation_parameters). We can also set the parameter called thinkingBudget, which defines the strategy for reasoning of this LLM. The openai package provides access to similar tuning parameters of gpt-5.1-chat-latest. In our experiments, we use the default values of all tuning parameters of gpt-5.1-chat-latest and gemini-2.5-flash, except for the parameter that controls their reasoning. For gpt-5.1-chat-latest, we set the reasoning-effort parameter to “medium.” For gemini-2.5-flash, we set thinkingBudget to “-1” and use the dynamic thinking strategy, which lets the LLM decide when and how much to think based on the complexity of the request.
Running an LLM using an API key has associated costs that depend on the number of tokens involved in the input and output of the model. Essentially, tokens are the data units of an LLM. They are constructed from text by the LLM using a process called tokenization, which considers language, punctuation, and context, among other features. For input and output text, gpt-5.1-chat-latest charges \$1.5 and \$10 for one million tokens, respectively (openai_gpt5_1_chat_latest). For gemini-2.5-flash, these costs are \$0.3 and \$2.5 (google_gemini_api_pricing_2025). Therefore, working with long prompts and outputs is more expensive than with short ones. In this regard, a benefit of our Zero-shot-CoT prompt template is that it allows us to limit the costs of our numerical experiments, because they involve many executions of Prompt 2.
@@ 4 Numerical results
We evaluate the compliant designs in terms of resolution and minimum moment aberration only, since they are computationally cheap and the latter is equivalent to minimum aberration. In Section 4.1, we show the numerical results of the GPT-5.1 model for the 8-, 16-, and 32-run design tasks separately. In Section 4.2, we discuss the performance of the Gemini 2.5 Flash model on these tasks.
@@ 4.1.1 Tasks with eight runs
Table 2 shows the performance of the GPT-5.1 model on the design construction tasks with eight runs. Specifically, the table shows the task label and the maximum, minimum, and median resolutions of the designs obtained by the LLM. For each task, the 10 designs constructed by the LLM were compliant.
@@
Table 2 shows that the resolution of the designs with four and seven factors is four and three, respectively. For five and six factors, the resolution is two or three. In contrast, all two-level fractional factorial designs with eight runs and four to seven factors in wu2011experiments have a resolution of at least three. The same is true for the 16- and 32-run fractional factorial designs, which we discuss later.
For each number of factors, the resolution of the best design obtained by the GPT-5.1 model equals that of the minimum aberration design in wu2011experiments. To further compare these designs, we computed their moment aberration patterns. Table 3 shows the patterns of the best designs obtained by the LLM for each task. The best designs obtained by this LLM have the same moment aberration patterns as the minimum aberration designs. Therefore, this LLM was able to construct the best designs for the 8-run design tasks. Regarding the consistency in obtaining them, the GPT-5.1 model found the best 4-, 5-, 6-, and 7-factor designs in 10, 8, 5, and 10, respectively, of the 10 independent executions of the corresponding Prompt 2; see Table 3.
Table 3: Moment aberration patterns obtained by the GPT-5.1 model for the 8-run design construction tasks.
@@ 4.1.2 Tasks with 16 runs
Table 4 shows the resolutions of the designs obtained by the GPT-5.1 model for the 16-run design tasks. The table also shows the number of compliant designs that were generated for each task. The LLM obtained 10 compliant designs for each task, except for the tasks with 12 to 15 factors. For these tasks, it only obtained eight or nine compliant designs. A close inspection to the non-compliant designs revealed that they had missing entries or were infeasible for the LLM. For the latter case, the LLM output a sentence similar to: “I’m sorry, but I cannot generate a valid design that meets your requirements.”
Table 4 shows that the median resolutions of the 16-run designs are higher than or equal to three, except for the designs with 9, 11, 14, and 15 factors. For these designs, the median resolution was 1.5, 2.5 or 2.
@@
For all numbers of factors, there were multiple compliant designs with the same resolution as the corresponding minimum aberration designs (wu2011experiments). Table 5 shows the frequencies of the moment aberration patterns of these designs for each task. To enhance its layout, the table shows the first four elements of the moment aberration pattern. For each number of factors, the GPT-5.1 model obtained the minimum aberration design at least twice, except for 12 factors for which it did not find the minimum aberration design. Interestingly, for tasks with five to eight factors, this LLM obtained the minimum aberration design in at least eight repetitions of these tasks. Therefore, the GPT-5.1 model performs well in constructing two-level fractional factorial designs with 16 runs and up to eight factors.
@@ 4.1.3 Tasks with 32 runs
Table 6 shows the resolution and number of complaint designs obtained by the GPT-5.1 model for the 32-run design tasks. For 11, 13, 16, and 19 to 26 factors, this LLM did not find a compliant design with a resolution higher than two. For this reason, these numbers of factors are omitted in the table. Table 6 shows that there were few compliant designs for tasks with more than eight factors. Similar to the 16-run design tasks, the 32-run noncompliant designs had missing entries or were infeasible for the LLM.
@@
For 6-9, 12, 17, and 18 factors, the maximum resolution in Table 6 matches that of the minimum aberration designs (wu2011experiments). For these numbers of factors, Table 7 shows the first six elements of the moment aberration patterns of the designs with the highest resolution obtained by the GPT-5.1 model. For 6-9 and 12 factors, the LLM found the minimum aberration design at least once. This is because the best moment aberration pattern is equal to that of the minimum aberration design for these cases. For 17 and 18 factors, the best moment aberration patterns are worse than those of the minimum aberration designs. Overall, we conclude that the GPT-5.1 model is moderately effective for constructing the 32-run 6-factor minimum aberration design, but ineffective for generating 32-run minimum aberration designs with more factors.
@@
4.2 Designs constructed using the Gemini-2.5 Flash model
We performed the design construction tasks with 8, 16, and 32 runs using the Gemini 2.5 Flash model. Remarkably, this LLM excelled in the 8-run design tasks because it obtained the minimum aberration designs for all factors in all independent executions of Prompt 2. This LLM also obtained 10 compliant designs for each 16-run design task. Table 8 shows the resolutions of these designs.
The Gemini 2.5 Flash model constructed 16-run designs with the same resolution as the minimum aberration designs. Table 9 summarizes the moment aberration patterns of the 16-run designs constructed by this LLM. The table is akin to Table 5 for the GPT-5.1 model, and it concerns the designs with the highest resolution obtained for each task. Except for 11 and 12 factors, the Gemini 2.5 Flash model obtained the 16-run minimum aberration designs at least twice; see Table 9.
@@
Regarding the 32-run design tasks, the LLM constructed 10 compliant designs for each number of factors, except for 17, 18, and 20 to 26 factors. For these cases, the LLM generated several non-compliant designs with missing entries or other issues. For tasks with 6-9, 10, and 14 factors, the compliant designs had a resolution higher than or equal to three. The distribution of the resolution is summarized in Table 10 for these cases. For six to nine factors, the maximum resolutions in this table equal those of comparable 32-run minimum aberration designs. Table 11, which is akin to Table 7, shows that the Gemini 2.5 Flash model obtained the minimum aberration designs with six to eight factors at least once. Remarkably, for six factors, the LLM constructed the 32-run minimum aberration design in the 10 independent executions of the corresponding Prompt 2.
@@
Table 11: Moment aberration patterns obtained by the Gemini-2.5 Flash model for selected 32-run design tasks. : Design has minimum moment aberration.
@@ 5 Conclusion
Based on our results, we recommend Gemini with its Gemini 2.5 Flash model to construct 8-run designs with four to seven factors, and a 32-run design with six factors using Prompt 2. We also recommend Gemini and this prompt to generate 16-run designs with five to eight factors, but warn practitioners that there is a small chance that this chatbot will produce poor designs. On the upside, all these numbers of factors are the most common in practical applications according to the systematic surveys of li2006regularities and ockuly2017response on real factorial experiments. On the downside, the Gemini 2.5 Flash model is infective in constructing attractive designs of other sizes consistently using Prompt 2. For these cases, we advise practitioners to use DoE textbooks (wu2011experiments; montgomery2017design), JMP, Minitab, or the FrF2 package in R.
@@
There are two potential ways to improve the performance of the GPT-5.1 and Gemini 2.5 Flash models on our design construction tasks. One way is to use a few-shot prompt template with CoT-based examples instead of our Zero-shot-CoT prompt template. Despite the improvement in performance of LLMs given by Zero-shot-CoT prompts, kojima2022large show that these prompts do not outperform few-shot prompts with CoT-based examples in general. However, developing these examples is challenging because their writing must have all the steps needed to reach the solution of a task’s instance. Moreover, the writing style of these steps influences the performance of LLMs (kojima2022large). In our setting, developing CoT-based examples would need a careful selection of two-level fractional factorial designs, as they can have different resolutions and WLPs. Moreover, each example would need a comprehensive step-by-step explanation of their construction. Such an explanation should include the selection of basic and generated factors, the elaboration of the defining relation and WLP, and their evaluation in terms of resolution and minimum aberration. For all these reasons, we leave the use of few-shot prompts with CoT-based examples for constructing designs as a topic for future research.
@@
Another way to improve the performance of the GPT-5.1 and Gemini 2.5 Flash models is using Retrieval-Augmented Generation (RAG; fan2024survey). RAG is a methodology for enhancing the performance of LLMs using domain-specific sources provided by the user. In our context, these sources can be textbooks and research articles on DoE that are open-source (Jones-Farmer19102024). RAG has three main steps. First, it stores the domain-specific sources in a database using encoding methods. Second, for an input prompt, it retrieves the elements from the database that are most similar to the prompt. Third, it informs the LLM about these elements by, for example, using them as the context of the prompt. In this way, RAG helps the LLM generate domain-specific output for a given task. We refer the reader to megahed2024introducing for an application of RAG in statistical quality control. We leave the application of RAG for constructing two-level fractional factorial designs as another topic for future research.
@@
Finally, we studied the default versions of the GPT and Gemini models used by ChatGPT’s and Gemini’s web interfaces at the time of writing this article (December, 2025). Due to the rapid development of research and applications of artificial intelligence, we expect the new versions of these LLMs—or even of other LLMs not studied here, such as the LLaMA model (touvron2023llama)—to become more capable in domain-specific tasks involving Zero-shot-CoT prompts. Our set of 36 design construction tasks thus enriches the catalog of benchmarks used to assess the gains in performance of these versions and LLMs (ni2025survey). We hope to see future studies showcasing the performance of these and other GenAI technologies in our set.

### v2

@@ 3.1.3 Zero-shot chain of thought
A chain of thought (CoT) is a series of intermediate natural language reasoning steps that can be added to a prompt to enhance the output of an LLM (Wei et al., 2022). This technique allows the LLM to decompose the main problem into a series of smaller intermediate problems, each solved before giving the final output. Wei et al. (2022) show that adding examples written using CoT to prompts improves the performance of transformer-based LLMs in arithmetic, symbolic reasoning, and common sense tasks.
Kojima et al. (2022) show that CoT can also improve the quality of zero-shot prompts. Specifically, they show that adding the sentence “Let’s think step by step,” or a similar text to a zero-shot prompt, generally enhances the performance of transformer-based LLMs in similar tasks as Wei et al. (2022). Kojima et al. (2022) refer to the technique of adding this sentence (or one of its variants) to a zero-shot prompt as Zero-shot-CoT. Attractive features of this technique include its simplicity, task independence, and the fact that it avoids the need for good CoT-based examples, which require significant expertise and careful writing.
@@ 3.1.4 Output format
We can set a format for an LLM to tailor the structure of their output to our needs (White et al., 2023). For example, we can provide constraints on the length and format of its output in the prompt. An output format is particularly useful for our problem because the LLMs tend to provide two-level designs with additional text to explain its properties and construction; see ChatGPT’s output 1 and Gemini’s output A1 in the Appendix. However, for our numerical experiments, we only need a table with the design in a specific format. That is, the table must have $m$ columns and $n$ rows with entries equal to 1-1 or 1. The table must also have a header row with the labels of the factor columns and an additional column to label the runs.
@@
The second element of our output format allows us to streamline the LLMs output because it sets the format for the table. Specifically, it instructs the LLM to generate the table using a comma-separated values (CSV) format, which is common for storing tables, matrices, and arrays. In this format, every entry in the table is separated by a comma. We specify the end of a row using two backslash symbols (\\). To set our second format element, we add the following sentences to our prompt: “The table must be in a comma-separated values (CSV) format. Specifically, the values in the table must be separated by ‘,’ and each row must end with ‘\\’. In the table, the first row will be used as a header row to label the factors using the letters of the English alphabet starting with ‘A’. The first column will be called “Run” and used to count the number of runs starting at ‘1’. Each design cell (excluding the header and Run columns) must contain either ‘-1’ or ‘1’.” We instruct the LLMs to label factors using the letters of the English alphabet following Montgomery (2017), ChatGPT’s output 1, and Gemini’s output A1. Moreover, the maximum number of factors in our experiments in Section 4 is 26, which is the number of letters in this alphabet.
@@ 3.2 Setup for the LLMs
To construct designs with Prompt 2, we use the GPT-5.1 and Gemini 2.5 Flash models, which are the default LLMs in ChatGPT’s and Gemini’s web interfaces, respectively, at the time of writing this article. Technically, these LLMs are gpt-5.1-chat-latest (OpenAI, 2025b) and gemini-2.5-flash (Google, 2025a). We run them using the Python packages called openai (OpenAI, 2023) and google-genai (Google, 2023), and API keys from OpenAI and Google. An API key is a unique code that can be used to access and communicate with an API, which, in turn, provides access to an LLM. The openai and google-genai packages provide user-friendly Python functions to interact with the APIs.
@@
The Python packages and APIs enable more tuning parameters of the LLMs than ChatGPT’s and Gemini’s web interfaces. For example, through the google-genai package, we can set the parameters called temperature, top-k, and top-p of gemini-2.5-flash, which control the randomness and diversity of its output to multiple runs of the same prompt (Google Cloud, 2025). We can also set the parameter called thinkingBudget, which defines the strategy for reasoning of this LLM. The openai package provides access to similar tuning parameters of gpt-5.1-chat-latest. In our experiments, we use the default values of all tuning parameters of gpt-5.1-chat-latest and gemini-2.5-flash, except for the parameter that controls their reasoning. For gpt-5.1-chat-latest, we set the reasoning-effort parameter to “medium.” For gemini-2.5-flash, we set thinkingBudget to “-1” and use the dynamic thinking strategy, which lets the LLM decide when and how much to think based on the complexity of the request.
Running an LLM using an API key has associated costs that depend on the number of tokens involved in the input and output of the model. Essentially, tokens are the data units of an LLM. They are constructed from text by the LLM using a process called tokenization, which considers language, punctuation, and context, among other features. For input and output text, gpt-5.1-chat-latest charges \$1.5 and \$10 for one million tokens, respectively (OpenAI, 2025b). For gemini-2.5-flash, these costs are \$0.3 and \$2.5 (Google AI for Developers, 2025). Therefore, working with long prompts and outputs is more expensive than with short ones. In this regard, a benefit of our Zero-shot-CoT prompt template is that it allows us to limit the costs of our numerical experiments, because they involve many executions of Prompt 2.
@@ 4 Numerical results
We evaluate the compliant designs in terms of resolution and minimum moment aberration calculated as in Section 2.2. In Section 4.1, we show the numerical results of the GPT-5.1 model for the 8-, 16-, and 32-run design tasks separately. In Section 4.2, we discuss the performance of the Gemini 2.5 Flash model on these tasks.
@@ 4.1.1 Tasks with eight runs
Table 2 shows the performance of the GPT-5.1 model on the design construction tasks with eight runs. Specifically, the table shows the task label and the maximum, minimum, and median resolutions of the designs obtained by this LLM. For each task, the 10 designs constructed by the LLM were compliant.
@@
Table 2 shows that the resolution of the designs with four and seven factors is four and three, respectively. For five and six factors, the resolution is two or three. In contrast, all two-level fractional factorial designs with eight runs and four to seven factors in Wu and Hamada (2011) have a resolution of at least three. The same is true for the 16- and 32-run fractional factorial designs, which we discuss later.
For each number of factors, the resolution of the best design obtained by the GPT-5.1 model equals that of the minimum aberration design in Wu and Hamada (2011). To further compare these designs, we computed their moment aberration patterns. Table 3 shows the patterns of the best designs obtained by the LLM for each task. The best designs have the same moment aberration pattern as the minimum aberration designs. Therefore, the GPT-5.1 model constructed the optimal designs for the 8-run design tasks. Regarding the consistency in obtaining them, it found the optimal 4-, 5-, 6-, and 7-factor designs in 10, 8, 5, and 10, respectively, of the 10 independent executions of Prompt 2; see Table 3.
Table 3: Moment aberration patterns obtained by the GPT-5.1 model for the 8-run design tasks.
@@ 4.1.2 Tasks with 16 runs
Table 4 summarizes the resolutions of the designs obtained by the GPT-5.1 model for the 16-run design tasks. The table also shows the number of compliant designs that were generated for each task. The LLM obtained 10 compliant designs for each task, except for the tasks with 12 to 15 factors. For these tasks, it only obtained eight or nine compliant designs. A close inspection to the non-compliant designs revealed that they had missing entries or were infeasible for the LLM. For the latter case, the LLM output a sentence similar to: “I’m sorry, but I cannot generate a valid design that meets your requirements.”
Table 4 shows that the median resolutions of the 16-run designs are higher than or equal to three, except for the designs with 11, 14, and 15 factors. For these designs, the median resolution was 1.5, 2 or 2.5. In general, the compliant designs of resolution one and two were not $2^{m-p}$ designs. This was also the case for the 32-run designs in the next section, and the designs obtained by the Gemini 2.5 Flash model in Section 4.2.
@@
For all numbers of factors, there were multiple compliant designs with the same resolution as the corresponding minimum aberration designs (Wu and Hamada, 2011). Table 5 shows the frequencies of the moment aberration patterns of these designs for each task. To enhance its layout, the table shows the first four elements of the moment aberration pattern. For each number of factors, the GPT-5.1 model obtained the minimum aberration design at least twice, except for 12 factors for which it did not find the minimum aberration design. Interestingly, for tasks with five to eight factors, this LLM obtained the minimum aberration design in at least eight repetitions of these tasks. Therefore, the GPT-5.1 model performs well in constructing the optimal two-level fractional factorial designs with 16 runs and up to eight factors.
@@ 4.1.3 Tasks with 32 runs
Table 6 shows the resolution and number of compliant designs obtained by the GPT-5.1 model for the 32-run design tasks. For 11, 13, 16, and 19 to 26 factors, this LLM did not find a compliant design with a resolution higher than two. For this reason, these numbers of factors are omitted in the table. Table 6 shows that there were few compliant designs for tasks with more than eight factors. Similar to the 16-run design tasks, the 32-run noncompliant designs had missing entries or were infeasible for the LLM.
@@
For 6-9, 12, 17, and 18 factors, the maximum resolution in Table 6 matches that of the minimum aberration designs (Wu and Hamada, 2011). For these numbers of factors, Table 7 shows the first six elements of the moment aberration pattern of the designs with the highest resolution obtained by the GPT-5.1 model. For 6-9 and 12 factors, the LLM found the minimum aberration design at least once. This is because the best moment aberration pattern is equal to that of the minimum aberration design for these cases. For 17 and 18 factors, the best moment aberration patterns are worse than those of the minimum aberration designs. Overall, we conclude that the GPT-5.1 model is moderately effective for constructing the 32-run 6-factor minimum aberration design, but ineffective for generating 32-run minimum aberration designs with more factors.
@@
4.2 Designs constructed using the Gemini 2.5 Flash model
We performed the design construction tasks with 8, 16, and 32 runs using the Gemini 2.5 Flash model. Remarkably, this LLM excelled in the 8-run design tasks because it obtained the minimum aberration designs for all factors in all independent executions of Prompt 2. This LLM also obtained 10 compliant designs for each 16-run design task. Table 8 summarizes the resolutions of these designs.
The Gemini 2.5 Flash model constructed 16-run designs with the same resolution as the minimum aberration designs. Table 9 shows the moment aberration pattern of the 16-run designs constructed by this LLM. The table is akin to Table 5 for the GPT-5.1 model, and it concerns the designs with the highest resolution obtained for each task. Except for 11 and 12 factors, the Gemini 2.5 Flash model obtained the 16-run minimum aberration designs at least twice; see Table 9.
@@
Regarding the 32-run design tasks, the LLM constructed 10 compliant designs for each number of factors, except for 17, 18, and 20 to 26 factors. For these cases, the LLM generated several non-compliant designs with missing entries or other issues. For tasks with 6-9, 10, and 14 factors, the compliant designs had a resolution higher than or equal to three. Table 10 shows the distribution of the resolution for these cases. For six to nine factors, the maximum resolutions in this table equal those of comparable 32-run minimum aberration designs. Table 11, which is akin to Table 7, shows that the Gemini 2.5 Flash model obtained the minimum aberration designs with six to eight factors at least once. Remarkably, for six factors, the LLM constructed the 32-run minimum aberration design in the 10 independent executions of the Prompt 2.
@@
Table 11: Moment aberration patterns obtained by the Gemini 2.5 Flash model for selected 32-run design tasks. : Design has minimum moment aberration.
@@ 5 Conclusion
Based on our results, we recommend Gemini with its Gemini 2.5 Flash model to construct 8-run designs with four to seven factors, and a 32-run design with six factors using Prompt 2. We also recommend Gemini and this prompt to generate 16-run designs with five to eight factors, but warn practitioners that there is a small chance that this chatbot will produce poor designs. On the upside, these numbers of factors are the most common in practice according to the systematic surveys of real factorial experiments conducted by Li et al. (2006) and Ockuly et al. (2017). On the downside, the Gemini 2.5 Flash model is infective in constructing attractive designs with more factors consistently using Prompt 2. For these cases, we advise practitioners to use DoE textbooks (Wu and Hamada, 2011; Montgomery, 2017), JMP, Minitab, or the FrF2 package in R.
@@
There are two potential ways to improve the performance of the GPT-5.1 and Gemini 2.5 Flash models on our design construction tasks. One way is to use a few-shot prompt template with CoT-based examples instead of our Zero-shot-CoT prompt template. Despite the improvement in performance of LLMs given by Zero-shot-CoT prompts, Kojima et al. (2022) show that these prompts do not outperform few-shot prompts with CoT-based examples in general. However, developing these examples is challenging because their writing must have all the steps needed to reach the solution of a task’s instance. Moreover, the writing style of these steps influences the performance of LLMs (Kojima et al., 2022). In our setting, developing CoT-based examples would need a careful selection of two-level fractional factorial designs, as they can have different resolutions and WLPs. Moreover, each example would need a comprehensive step-by-step explanation of their construction. Such an explanation should include the selection of basic and generated factors, the elaboration of the defining relation and WLP, and the evaluation in terms of resolution and minimum aberration. For all these reasons, we leave the use of few-shot prompts with CoT-based examples for constructing designs as a topic for future research.
@@
Another way to improve the performance of the GPT-5.1 and Gemini 2.5 Flash models is using Retrieval-Augmented Generation (RAG; Fan et al., 2024). RAG is a methodology for enhancing the performance of pre-trained LLMs using domain-specific sources provided by the user. In our context, these sources can be textbooks and research articles on DoE that are open-source (Jones-Farmer et al., 2024). RAG has three main steps. First, it stores the domain-specific sources in a database using encoding methods. Second, for an input prompt, it retrieves the elements from the database that are most similar to the prompt. Third, it informs the LLM about these elements by, for example, using them context in the prompt. In this way, RAG helps the LLM generate domain-specific output for a given task. We refer the reader to Megahed et al. (2024) for an application of RAG in statistical quality control, and leave its application for constructing two-level fractional factorial designs as another topic for future research.
@@
Finally, we studied the default versions of the GPT and Gemini models used by ChatGPT’s and Gemini’s web interfaces at the time of writing this article (December, 2025). Due to the rapid development of research and applications of artificial intelligence, we expect the new versions of these LLMs—or even of other LLMs not studied here, such as the LLaMA model (Touvron et al., 2023)—to become more capable in domain-specific tasks involving Zero-shot-CoT prompts. Our set of 36 design construction tasks thus enriches the catalog of benchmarks used to assess the gains in performance of these versions and LLMs (Ni et al., 2025). We hope to see future studies showcasing the performance of these and other GenAI technologies in our set.